Differential Equation
Q: Find the general solution to the given homogeneous problem.

10 a.) y + y - 2y - 2y = 0
b.) y(4) + 4y + 4y = 0
Answer #1

y+ y-2y-29=0 Auniliry equitim is fto%27-2=0 =-1, -Sz, Sz Generd sAubion is seit yL4)+44+ 4y =0 (6) Aunilisy eguahin is 42=